OpenStudy (anonymous):

somebody please I will fail help 9th grade algebra medal and fan!!!

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

1) For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (hint: you need to find the “k” value) In order for y to vary directly with x, it must satisfy the equation y = kx. So, for the values, plug them in to see: 6= k(9) 8 = k(10) 10= k(11 X Y 6 9 8 10 10 11 2) For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (Hint: you need to find the “k” value) X Y 8 6 12 9 20 15 3) Graph -1/6 x-1/9 y=-1/3 4) Graph x=4 (Hint: put the equation in slope-intercept form) 5) The table shows the number of miles driven over time. Express the relationship between distance and time in simplified form as a unit rate: ___________ mph. Write a statement that interprets the relationship: _________________________________________________. Time (hours) 3 5 7 9 Distance (miles) 183 305 427 549 6) Find the slope of the line. M= __________ 7) What is the slope of the line? M= _____________ Hint: slope is rise/run or (change in y)/(change in x) 8) What is the slope of a line passing through the pair of points (-5, -6) and (3, -1)? Hint: slope formula is (y_2-y_1)/(x_2-x_1 ) 9) What is the slope of a line passing through the pair of points (3/2,-10) and (3, 7/2)? Hint: slope formula is (y_2-y_1)/(x_2-x_1 ) 10) Write an equation in point-slope form for the line through the given point with the given slope. (3,2):m=4 Hint: Point-slope form is y-y_1=m(〖x-x〗_1) 11) The table shows the height of a plant as it grows. Write an equation in point-slope form that represents the plant’s height at any time. Determine the height of the plant after 1 year. Time Height 2 8 4 15 6 22 8 29 12) Write each equation in standard form. (Hint: Standard form has three requirements) a) y=-4/7 x+2 b) y=-0.3x+0.8

OpenStudy (anonymous):

plz help @AFlyingMelon I promise in is not a test it is half of my study guide and if I fail it I will fail it is worth 30% of my grade plz

OpenStudy (anonymous):

I do not know too much, but here is what I know: 1. Y does not vary directly with X. 2. Y is always 75% of X, so:\[Y = 0.75 \times X\]or\[Y = \frac{ 3}{ 4} \times X\]K would be 0.75(or 3/4), so the equation would be: \[Y = KX\]or\[Y = K \times X\] 5. Since Speed equals Distance divided by Time(S = D/T), the vehicle is traveling at 61 mph. The unit rate would be: \[\frac{ 61 }{ 1 }\]The statement for the unit rate would be "For every 1 hour the vehicle travels, the vehicle will travel 61 miles." I hope this helped, but maybe @thomaster will be able to help you.